# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v1

@@ 2.4 Necessity of IVX Transformation
$\widehat{\psi}^{(j)}$ $=\arg\min_{\psi}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{w}_{j,t-1}-W_{-j,t-1}^{\top}\psi)^{2}+\mu\|D_{-j}\psi\|_{1}.$ (2.15)
@@
where $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{p})^{\top}$ is a $p$-dimensional Brownian motion, $\sigma_{j}^{*}=\sqrt{\int_{0}^{1}\mathcal{B}_{j}^{2}-(\int_{0}^{1}\mathcal{B}_{j})^{2}}$, and $\mathcal{B}_{-j}=(\mathcal{B}_{k})_{k\neq j}$. In other words, the least square estimator converges in distribution to a nondegenerate stable law. This is the well-known spurious regression phenomenon in unit root regressions. While (2.16) suggests that the limit target coefficients are random and continuously distributed, the asymptotics for the LASSO estimator (2.15) in high dimension depends on the sparsity of the target coefficients. The randomness of coefficients is contradictory to the sparsity required in LASSO.
Second, as a consequence of the spurious regression, the least squares residual $\widehat{r}_{j,t}^{{\rm OLS}}=\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{(j,{\rm OLS})}$, which under large $n$ is expected to be approximated by $\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{*}$ in view of (2.16), remains highly persistent. Using $\widehat{r}_{j,t}^{{\rm OLS}}$ as the score vector would therefore keep the Stambaugh bias in the noise component ${\rm N}_{j}$. The same issue is also anticipated in the score vector $r_{j,t}$ in (2.14) by LASSO.
@@ 3 Asymptotic Theory
The following Assumption 2 imposes restrictions on the $\alpha$-mixing coefficients that characterize the time dependence of the innovation $\varepsilon_{t}$. For any two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$, define $\alpha(\mathcal{A},\mathcal{B})=\sup_{A\in\mathcal{A},B\in\mathcal{B}}|\Pr\left(A\cap B\right)-\Pr(A)\Pr(B)|$ and $\alpha(d)=\sup_{s\in\mathbb{Z}}\alpha(\sigma(\{u_{t},\varepsilon_{t}\}_{t\leq s}),\sigma(\{u_{t},\varepsilon_{t}\}_{t\geq s+d}))$.
@@
There exists some absolute constants $C_{\alpha}$, $c_{\alpha}$, $r$, $c_{\varepsilon}$ such that
@@
(a) $p=O(n^{\nu})$ for an arbitrary $\nu>0$ and (b) $s=O(n^{\frac{1}{4}(\tau\wedge(1-\tau))-\xi}\vee p^{1-\xi})$ for an arbitrary small $\xi>0$.
@@
In empirical finance, the m.d.s assumption is commonly imposed on the error term, especially for testing asset return predictability (Zhu et al., , 2014; Kostakis et al., , 2015). It indicates that the dependent variable (financial asset return) is not predictable if the null hypothesis of zero regression coefficients is not rejected, which aligns with the efficient market hypothesis. In macroeconomic applications, the high dimensional covariates with different degrees of persistence alleviate the concern of variable omission, which makes the m.d.s  assumption plausible. [^3] Mild model misspecification with approximate sparsity can be accommodated by our framework; see Belloni et al., (2012) and Mei and Shi, (2024, Remark 1).
@@
Third, Assumption 4(a) follows Mei and Shi, (2024) that allows the variable dimension $p$ to diverge at a polynomial rate of the sample size $n$; it can be extended to an exponential rate of $n$ at the cost of expositional complications. Nevertheless, Assumption 4(b) imposes a more restrictive condition for the sparsity index $n$, compared to $s=o(n^{1/4})$ in Mei and Shi, (2024), which is reasonable as asymptotic normality is more delicate and demanding than consistency. This condition ensures that the shrinkage bias is accurately estimated to achieve the result (R2) in Section 2.2, so that XDlasso asymptotically follows a normal distribution centered at the true coefficient; also see the second term of (3.13) below.
@@
We recommend $\tau=1/2$ for practitioners, under which the quantity $(\tau\wedge(1-\tau))/4$ achieves its maximum $1/8$ and thus permits the weakest sparsity condition. This is different from the conventional wisdom of IVX (Phillips and Lee, , 2016; Kostakis et al., , 2015) where $\tau$ is recommended to be as large as 0.95 to minimize the loss of rate efficiency (or local power). The context is different in our setting of high dimensional predictive regression, under which we require a very small sparsity index $s$ relative to $n$. An excessively large $\tau$ will thus cause a severe distortion in the size of the test under finite sample. As discussed in Theorem 2 below, the hypothesis testing based on XDlasso is consistent for a wide class of local-to-zero $\theta_{j}^{*}$, despite a slower convergence rate compared with the case when $\tau$ is close to 1.
@@
Under Assumptions 14, there exists some absolute constant $C_{\lambda}$ such that when the LASSO tuning parameter in (2.2) satisfies $\lambda=C_{\lambda}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$, as $n\to\infty$, we have
@@ 3.1 The Auxiliary Regression
While the convergence of Slasso in the main regression is well expected, new theory must be developed for the auxiliary regression (2.10), again by the workhorse Slasso when $w_{j,t}$ is either a unit root or stationary regressor. Importantly, as we maintain an agnostic attitude about the persistence of the variable of interest, which is the dependent variable of the auxiliary regression, the theory must be simultaneously applicable to the two very different types of time series in a unified manner.
Without loss of generality, let $\mathcal{M}_{x}=[p_{x}]$ and $\mathcal{M}_{z}=[p]\backslash[p_{x}]$ denote the integer sets indexing the locations of unit root and stationary regressors, respectively. Suppose $j\in\mathcal{M}_{x}.$ The standardized instrument $\widetilde{\zeta}_{j,t}$ has a different degree of persistence from $X_{-j,t}$ and $Z_{t}$, due to the different orders of integration. In the low dimensional framework, the FCLT in Phillips and Lee, (2016) yields the following rate of convergence in OLS:
@@
In high dimensional models, the sample Gram matrix is rank-deficient and the FCLT no longer works. Thanks to the $L_{1}$ penalization, Slasso has a comparable local-to-zero order as (3.7).
Proposition 1.
Suppose that $j\in\mathcal{M}_{x}$ and Assumptions 14 hold. There exists an absolute constant $C_{\mu}$ such that when $\mu=C_{\mu}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{\tau\wedge(1-\tau)}}$, we have
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}=O_{p}\left(\dfrac{(\log p)^{1+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right).$ (3.8)
Remark 6.
In Proposition 1, the specified rate for the tuning parameter $\mu$ closely links to the limit theory in low dimensions. Suppose the $(p-1)$-dimensional vector $\varphi^{(j)}$ is the target parameter of regression (2.10), such that $\|D_{-j}\varphi^{(j)}\|_{1}=O_{p}\left(1/\sqrt{n^{\tau\wedge(1-\tau)}}\right)$. Parallel to the FCLT in the low dimensional model, the cross-product between the standardized regressors $D_{-j}^{-1}W_{-j,t}$ and the error term of the regression $\widetilde{\zeta}_{j,t}-W_{-j,t}\varphi^{(j)}$ is bounded by
$\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t}\left(\widetilde{\zeta}_{j,t}-W_{-j,t}\varphi^{(j)}\right)\|_{\infty}=O_{p}\left(1/\sqrt{n^{\tau\wedge(1-\tau)}}\right).$
The convergence rate of the LASSO tuning parameter for high dimensional regression in Proposition 1 equals the rate of the low dimensional deviation bound $1/\sqrt{n^{\tau\wedge(1-\tau)}}$ multiplied by a logarithmic factor. This fact echoes the results for i.i.d. data (Bickel et al., , 2009) and high dimensional nonstationary regressors (Mei and Shi, , 2024).
When $j\in\mathcal{M}_{z}$, we abuse the notation $Z_{-j,t}$ to denote the vector of stationary regressors excluding $w_{j,t}$. Define
@@
In addition, the error term $\widetilde{\eta}_{j,t}$ is close to the stationary time series $(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$. Thus, the error term $\widetilde{\eta}_{j,t}$ is asymptotically uncorrelated to the nonstationary regressors $X_{t}$ due to different persistence. Furthermore, the coefficients $\varphi_{0z}^{(j)*}$ satisfy $\mathbb{E}\left(Z_{-j,t}(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})\right)=0$. Therefore, $\widetilde{\eta}_{j,t}$ is also asymptotically uncorrelated to the the stationary regressors $Z_{-j,t}$, thereby ensuring the consistency of the LASSO estimator $\widehat{\varphi}^{(j)}$ in (2.10).
@@
To bound the LASSO estimation errors, we need sparsity of not only the main regression (2.1), but also the auxiliary regression (3.11). This is very similar to the high dimensional sparse instrumental variable regression; see Zhu, (2018) and Gold et al., (2020). In Assumption 5, we slightly abuse the sparsity index $s$ to bound the number of nonzero coefficients in the vector $\varphi_{0z}^{(j)*}$. For simplicity, we directly impose this high-level sparsity assumption on $\varphi_{0z}^{(j)*}$, which can be deduced under the commonly used sparsity restriction on the precision matrix $\left(\mathbb{E}\left(Z_{t}Z_{t}^{\top}\right)\right)^{-1}$; see the definition of $s_{j}$ and the conditions in Theorem 2.1 of Zhang and Cheng, (2017, pp. 759). Finally, the upper bound of the $L_{1}$-norm of $\varphi_{0z}^{(j)*}$ controls the variance of the error term in the pseudo-true model (3.11).
The following proposition formally lays out the convergence rate of the auxiliary LASSO estimator (2.10) when $w_{j,t}$ is stationary.
@@
Suppose that $j\in\mathcal{M}_{z}$ and Assumptions 15 hold. Then, there exists some absolute constant $C_{\mu}>0$ such that when $\mu=C_{\mu}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, we have
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right)$
where $\varphi^{(j)*}=(0_{p_{x}}^{\top},\varphi_{z}^{(j)*\top})^{\top}$.
@@
The convergence rate of tuning parameter $\mu$ specified in Proposition 2 induces the following Karush–Kuhn–Tucker (KKT) condition
$\|n^{-1}\sum_{t=1}^{n}D_{-j}W_{-j,t}\widehat{r}_{j,t}\|_{\infty}\leq\dfrac{\mu}{2}\leq\dfrac{C_{\mu}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (3.12)
which is sufficient to bound the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$ in (2.5). Nevertheless, we still need the consistency of $\widehat{\varphi}_{j}$ established in Propositions 1 and 2. These consistency results not only help us show the asymptotic normality of the $t$-statistic $t_{j}^{{\rm XD}}$ defined in (2.13), which guarantees the asymptotic size of the test, but are also necessary for establishing the convergence rate of the standard error that governs the asymptotic power.
@@ 3.2 Asymptotic Normality
The consistency of the auxiliary regression built in Propositions 1 and 2 guarantees the following decomposition of the statistic:
@@
where ${\rm sgn}_{j}=\frac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is either 1 or $-1$ with probability one as $\frac{\left|\sum_{t=1}^{n}\widetilde{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is continuously distributed. The first term of the above expression is a counterpart of $\mathrm{N}_{j}/\omega_{j}$ in the discussion of generic desparsifying argument in (2.5). The second term is the convergence rate of the approximation error of the shrinkage bias analogous to $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}$. Under Assumption 4, the second term of the right-hand side of (3.13) is asymptotically negligible, thereby yielding the following asymptotic normality for any $j\in[p]$.
@@
Suppose Assumptions 1-5 hold. There exist absolute constants $C_{\lambda}$ and $C_{\mu}$ such that when $\lambda=C_{\lambda}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\mu=C_{\mu}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, respectively, as $n\to\infty$, we have
$\left(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}\right)\big{/}\widehat{\omega}_{j}^{{\rm XD}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (3.14)
@@
The convergence rates displayed in Theorem 2 are coherent with the results in low dimensional predictive regressions. When $j\in\mathcal{M}_{x}$, the standard error converges faster than the parametric rate $1/n^{\delta_{j}}$ for any $\delta_{j}\in(0,(1+\tau)/2)$. Thus, the hypothesis testing based on the $t$-statistic $t_{j}^{{\rm XD}}$ is consistent under the alternative with $\beta_{j}^{*}=c/n^{\delta_{j}}$ for some $c>0$ over this class of $\delta_{j}$. The range $\delta_{j}\in(0,(1+\tau)/2)$ includes the important $1/\sqrt{n}$ rate of coefficients for a unit root $X_{j,t}$, under which ${\rm var}(X_{j,t}\beta_{j}^{*})=O(1).$ The $1/\sqrt{n}$ factor thus balances the larger order of unit root regressors $X_{t}$ and the standard $O_{p}(1)$ stochastic order of $y_{t}$ (Phillips, , 2015; Lee et al., , 2022), and XDlasso is consistent under this class of alternatives. When $j\in\mathcal{M}_{z}$, XDlasso achieves the standard $\sqrt{n}$-consistency.

### v2

@@ 2.4 Necessity of IVX Transformation
$\widehat{\psi}^{(j)}$ $=\arg\min_{\psi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{w}_{j,t-1}-W_{-j,t-1}^{\top}\psi)^{2}+\mu\|D_{-j}\psi\|_{1}.$ (2.15)
@@
where $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{p})^{\top}$ is a $p$-dimensional Brownian motion, $\sigma_{j}^{*}=\sqrt{\int_{0}^{1}\mathcal{B}_{j}^{2}-(\int_{0}^{1}\mathcal{B}_{j})^{2}}$, and $\mathcal{B}_{-j}=(\mathcal{B}_{k})_{k\neq j}$. In other words, the least square estimator converges in distribution to a nondegenerate random variable. This is the well-known spurious regression phenomenon in unit root regressions. While (2.16) suggests that the limit target coefficients are random and continuously distributed, the asymptotics for the LASSO estimator (2.15) in high dimensions depends on the sparsity of the target coefficients. The randomness of coefficients is contradictory to the sparsity required in LASSO.
Second, as a consequence of the spurious regression, the least squares residual $\widehat{r}_{j,t}^{{\rm OLS}}=\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{(j,{\rm OLS})}$, which under large $n$ is expected to be approximated by $\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{*}$ in view of (2.16), remains highly persistent. Using $\widehat{r}_{j,t}^{{\rm OLS}}$ as the score vector in low dimensions would therefore keep the Stambaugh bias in the noise component ${\rm N}_{j}$. The same issue is also anticipated in the score vector $r_{j,t}$ in (2.14) by LASSO in high dimensions.
@@ 3 Asymptotic Theory
The following Assumption 2 imposes restrictions on the $\alpha$-mixing coefficients that characterize the time dependence of the innovations $u_{t}$ and $\varepsilon_{t}$. For any two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$, define $\alpha(\mathcal{A},\mathcal{B})=\sup_{A\in\mathcal{A},B\in\mathcal{B}}|\Pr\left(A\cap B\right)-\Pr(A)\Pr(B)|$ and $\alpha(d)=\sup_{s\in\mathbb{Z}}\alpha(\sigma(\{u_{t},\varepsilon_{t}\}_{t\leq s}),\sigma(\{u_{t},\varepsilon_{t}\}_{t\geq s+d}))$.
@@
There exist some absolute constants $C_{\alpha}$, $c_{\alpha}$, $r$, $c_{\varepsilon}$ such that
@@
(a) $p=O(n^{\nu})$ for an arbitrary $\nu>0$ and (b) $s=O(n^{\frac{1}{4}(\tau\wedge(1-\tau))-\xi}\wedge p^{1-\xi})$ for an arbitrary small $\xi>0$.
@@
In empirical finance, the m.d.s assumption is commonly imposed on the error term, especially for testing asset return predictability (Zhu et al.,, 2014; Kostakis et al.,, 2015). It indicates that the dependent variable (financial asset return) is not predictable if the null hypothesis of zero regression coefficients is not rejected, which aligns with the efficient market hypothesis. In macroeconomic applications, the high dimensional covariates with different degrees of persistence alleviate the concern of variable omission, which makes the m.d.s. assumption plausible. [^3] Mild model misspecification with approximate sparsity can be accommodated by our framework; see Belloni et al., (2012) and Mei and Shi, (2024, Remark 1).
@@
Third, Assumption 4(a) follows Mei and Shi, (2024) that allows the variable dimension $p$ to diverge at a polynomial rate of the sample size $n$; it can be extended to an exponential rate of $n$ at the cost of expositional complications. Assumption 4(b) imposes a more restrictive condition for the sparsity index $s$, compared to $s=o(n^{1/4})$ in Mei and Shi, (2024), which is reasonable as asymptotic normality is more delicate and demanding than consistency. This condition ensures that the shrinkage bias is accurately estimated to achieve the result (R2) in Section 2.2, so that XDlasso asymptotically follows a normal distribution centered at the true coefficient; also see the second term of (3.13) below.
@@
In practice the sparsity index $s$ is unknown. We therefore recommend $\tau=1/2$ for practitioners, under which the quantity $(\tau\wedge(1-\tau))/4$ achieves its maximum $1/8$ and thus permits the weakest sparsity condition. This is different from the conventional wisdom of IVX (Phillips and Lee,, 2016; Kostakis et al.,, 2015) where $\tau$ is recommended to be as large as 0.95 to minimize the loss of rate efficiency (or local power). The context is different in our setting of high dimensional predictive regression, under which An excessively large $\tau$ will require a very small sparsity index $s$ relative to $n$ and thus cause a severe distortion in the size of the test under finite sample. As discussed in Theorem 2 below, the hypothesis testing based on XDlasso is consistent for a wide class of local-to-zero $\theta_{j}^{*}$, despite a slower convergence rate compared with the case when $\tau$ is close to 1.
@@
Under Assumptions 14, there exists some absolute constant $C_{{\rm m}}$ such that when the LASSO tuning parameter (2.2) for the main regression satisfies $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$, as $n\to\infty$, we have
@@ 3.1 The Auxiliary Regression
While the convergence of Slasso in the main regression is well expected, new theory must be developed for the auxiliary regression (2.10), again by the workhorse Slasso when $w_{j,t}$ is either a unit root or stationary regressor. Importantly, as we maintain an agnostic attitude about the persistence of the variable of interest, which is the dependent variable of the auxiliary regression, the theory must be simultaneously applicable to the two very different types of time series in a unified manner. Without loss of generality, let $\mathcal{M}_{x}=[p_{x}]$ and $\mathcal{M}_{z}=[p]\backslash[p_{x}]$ denote the integer sets indexing the locations of unit root and stationary regressors, respectively.
We first examine the case when $j\in\mathcal{M}_{x}.$ The standardized instrument $\widetilde{\zeta}_{j,t}$ has a different degree of persistence from $X_{-j,t}$ and $Z_{t}$, due to the different orders of integration. In the low dimensional framework, the FCLT in Phillips and Lee, (2016) yields the following rate of convergence in OLS:
@@
In high dimensional models, the sample Gram matrix is rank-deficient and the FCLT no longer works. Thanks to the $L_{1}$ penalization, Slasso has a comparable local-to-zero order as (3.7), which is shown in Proposition 1 below.
We then turn to $j\in\mathcal{M}_{z}$, and slightly abuse the notation $Z_{-j,t}$ to denote the vector of stationary regressors excluding $w_{j,t}$. Define
@@
In addition, the error term $\widetilde{\eta}_{j,t}$ is close to the stationary time series $(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$ and thus asymptotically uncorrelated to the nonstationary regressors $X_{t}$ due to different persistence. Furthermore, the coefficients $\varphi_{0z}^{(j)*}$ satisfy $\mathbb{E}\left(Z_{-j,t}(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})\right)=0$. Therefore, $\widetilde{\eta}_{j,t}$ is also asymptotically uncorrelated to the the stationary regressors $Z_{-j,t}$, thereby ensuring the consistency of the LASSO estimator $\widehat{\varphi}^{(j)}$ in (2.10).
@@
To bound the LASSO estimation errors, we need sparsity of not only the main regression (2.1), but also the auxiliary regression (3.10). This is similar to the high dimensional sparse instrumental variable regression; see Zhu, (2018) and Gold et al., (2020). In Assumption 5, we slightly abuse the sparsity index $s$ to bound the number of nonzero coefficients in the vector $\varphi_{0z}^{(j)*}$. For simplicity, we directly impose this high-level sparsity assumption on $\varphi_{0z}^{(j)*}$, which can be deduced under the commonly used sparsity restriction on the precision matrix $\left(\mathbb{E}\left(Z_{t}Z_{t}^{\top}\right)\right)^{-1}$; see the definition of $s_{j}$ and the conditions in Theorem 2.1 of Zhang and Cheng, (2017, pp. 759). Finally, the upper bound of the $L_{1}$-norm of $\varphi_{0z}^{(j)*}$ controls the variance of the error term in the pseudo-true model (3.10). Parallel conditions naturally hold for $j\in\mathcal{M}_{x}$ as the pseudo-true coefficients are zero in view of the local-to-zero OLS estimate displayed in (3.7).
The following proposition formally lays out the convergence rate of the auxiliary LASSO estimator (2.10).
@@
Suppose that $j\in\mathcal{M}_{z}$ and Assumptions 15 hold. Then, there exists some absolute constant $C_{{\rm a}}>0$ such that when $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, we have
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right)$ (3.11)
where $\varphi^{(j)*}={\bf 1}\{j\in\mathcal{M}_{x}\}\cdot 0_{p-1}+{\bf 1}\{j\in\mathcal{M}_{z}\}\cdot(0_{p_{x}}^{\top},\varphi_{z}^{(j)*\top})^{\top}$.
@@
The rate of tuning parameter $\mu$ specified in Proposition 1 induces the following Karush–Kuhn–Tucker (KKT) condition
$\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t}\widehat{r}_{j,t}\|_{\infty}\leq\dfrac{\mu}{2}\leq\dfrac{C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (3.12)
which is sufficient to bound the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$ in (2.5). Nevertheless, it is still necessary to establish the consistency of $\widehat{\varphi}^{(j)}$. This consistency result not only helps us show the asymptotic normality of the $t$-statistic $t_{j}^{{\rm XD}}$ defined in (2.13), which guarantees the asymptotic size of the test, but is also critical for the convergence rate of the standard error that governs the asymptotic power.
@@ 3.2 Asymptotic Normality
The consistency of the auxiliary regression built in Proposition 1 guarantees the following decomposition of the statistic:
@@
where ${\rm sgn}_{j}=\frac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is either 1 or $-1$ with probability one as $\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}$ is continuously distributed. The first term of the above expression is a counterpart of $\mathrm{N}_{j}/\omega_{j}$ in the discussion of generic desparsifying argument in (2.5). The second term is the convergence rate of the approximation error of the shrinkage bias analogous to $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}$. Under Assumption 4, the second term of the right-hand side of (3.13) is asymptotically negligible, thereby yielding the following asymptotic normality for any $j\in[p]$.
@@
Suppose Assumptions 1-5 hold. There exist absolute constants $C_{{\rm m}}$ and $C_{{\rm a}}$ such that when $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, respectively, as $n\to\infty$, we have
$(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*})\big{/}\widehat{\omega}_{j}^{{\rm XD}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (3.14)
@@
The convergence rates displayed in Theorem 2 are coherent with the results in low dimensional predictive regressions. When $j\in\mathcal{M}_{x}$, the standard error converges faster than the rate $1/n^{\delta_{j}}$ for any $\delta_{j}\in(0,(1+\tau)/2)$. Thus, for the null hypothesis $\mathbb{H}_{0}:\beta_{j}^{*}=0$, the hypothesis testing based on the $t$-statistic $t_{j}^{{\rm XD}}$ is consistent under the alternative with $\beta_{j}^{*}=c/n^{\delta_{j}}$ for some $c>0$ over this class of $\delta_{j}$. The range $\delta_{j}\in(0,(1+\tau)/2)$ includes the important $1/\sqrt{n}$ rate of coefficients for a unit root $X_{j,t}$, under which ${\rm var}(X_{j,t}\beta_{j}^{*})=O(1).$ The $1/\sqrt{n}$ factor thus balances the larger order of unit root regressors $X_{t}$ and the standard $O_{p}(1)$ stochastic order of $y_{t}$ (Phillips,, 2015; Lee et al.,, 2022), and the test by XDlasso is consistent under this class of alternatives. When $j\in\mathcal{M}_{z}$, XDlasso achieves the standard $\sqrt{n}$-consistency.
Remark 7.
Unlike the convergence rates of Slasso in Lemma 1, the convergence rate of XDlasso in Theorem 2 is independent of the variable dimension $p$ and the sparsity index $s$. In view of (2.5), in XDlasso the order of the approximation error of the shrinkage bias, $\widehat{{\rm B}}_{j}-{\rm B}_{j}$, depends on $p$ and $s$, but it is dominated by the order of the noise component ${\rm N}_{j}$. The convergence rate of XDlasso therefore follows the order of ${\rm N}_{j}$, which relates only to the sample size $n$.